PLEASE EXPLAIN HOW TO SOLVE IT WITH EXCEL

1)

A researcher wanted to determine which method would help smokers quit smoking. He divided a group of 183 smokers into three groups: Group 1 is given a patch, Group 2 is given oral medication, and Group 3 has no intervention. The results of the study are in the below table.

Group 1 Group 2 Group 3
Did Not Quit 52 54 51
Quit 13 9 4

*Source: Harris Poll

  1. Determine the null and alternative hypotheses.
    • H0:H0: The methods for helping people quit smoking are the same.
      Ha:Ha: The methods for helping people quit smoking follow a different distribution.
    • H0:H0:The methods for helping people quit smoking are independent.
      Ha:Ha: The methods for helping people quit smoking are dependent.

  2. Determine the test statistic. Round your answer to two decimal places.
    χ2 = Incorrect3.96
  3. Determine the p-value. Round your answer to four decimal places.
    p-value =
  4. Make a decision.
    • Fail to reject the null hypothesis.
    • Reject the null hypothesis.

  5. Make a conclusion.
    • There is sufficient evidence to support the claim that the methods for helping people quit smoking are different.
    • There is not sufficient evidence to support the claim that the methods for helping people quit smoking are different.

2)

Pain medications have side effects where one common side effect is dizziness. A researcher wanted to determine whether some pain medications produce more or less dizziness than others. The results from his study are shown below in the table. Apply a test for homogeneity to determine whether the proportion of people taking a pain medication within each treatment group that experience dizziness are the same at the α=0.01α=0.01 level of significance.

Tordol Placebo Naproxen Tylenol Advil
Dizziness 83 32 36 5 8
No dizziness 4,063 1,832 1,330 382 337
  1. Determine the null and alternative hypotheses.
    • H0:H0:The proportion of people within each treatment group who experienced dizziness are independent.
      Ha:Ha: The proportion of people within each treatment group who experienced dizziness are dependent.
    • H0:H0: The proportion of people within each treatment group who experienced dizziness are the same.
      Ha:Ha: The proportion of people within each treatment group who experienced dizziness are different.

  2. Determine the test statistic. Round your answer to two decimal places.
    χ2=χ2=
  3. Determine the pp-value. Round your answer to four decimal places.
    pp-value =
  4. Make a decision.
    • Fail to reject the null hypothesis.
    • Reject the null hypothesis.

  5. Make a conclusion.
    • There is not sufficient evidence to support the claim that the proportion of people within each treatment group who experienced dizziness are different.
    • There is sufficient evidence to support the claim that proportion of people within each treatment group who experienced dizziness are different.
Answer #1

1.

H0:The methods for helping people quit smoking are independent.
Ha: The methods for helping people quit smoking are dependent.

Observed frequencies:

Group 1 Group 2 Group 3 Totals
Did not quit 52=O11 54=O12 51=O13 157=O10
Quit 13=O21 9=O22 4=O23 26=O20
Totals 65=O01 63=O02 55=O03 183=N

Expected frequencies:

E_{ij}=\frac{O_{i0}O_{0j}}{N}

Group 1 Group 2 Group 3
Did not quit 55.765=E11 54.0492=E12 47.1858=E13
Quit 9.235=E21 8.9508=E22 7.8142=E23

Following~table~shows~\frac{(O_{ij}-E_{ij})^2}{E_{ij}}:

0.2542 0 0.3083
1.5349 0.0003 1.8618

Value~of~\chi^2=\sum_{i=1}^2\sum_{j=1}^3\frac{(O_{ij}-E_{ij})^2}{E_{ij}}=3.96\\\\ P-value=P(\chi^2>3.96|\chi^2\sim \chi^2_2)=0.1381\\\\ (R~code:~round(1-pchisq(3.96,2),4))

Since P-value>0.05, Option: Fail to reject the null hypothesis.

Conclusion: There is not sufficient evidence to support the claim that the methods for helping people quit smoking are different.

2.

H0: The proportion of people within each treatment group who experienced dizziness are the same.
Ha: The proportion of people within each treatment group who experienced dizziness are different.

Observed frequencies:

Tordol Placebo Naproxen Tylenol Advil Totals
Did not quit 83=O11 32=O12 36=O13 5=O14 8=O15 164=O10
Quit 4,063=O21 1,832=O22 1,330=O23 382=O24 337=O25 7,944=O20
Totals 4146=O01 1864=O02 1366=O03 387=O04 345=O05 8108=N

Expected frequencies:

E_{ij}=\frac{O_{i0}O_{0j}}{N}

Tordol Placebo Naproxen Tylenol Advil
Did not quit 83.8609=E11 37.703=E12 27.63=E13 7.8278=E14 6.9783=E15
Quit 4062.139=E21 1826.297=E22 1338.37=E23 379.1722=E24 338.0217=E25

Following~table~shows~\frac{(O_{ij}-E_{ij})^2}{E_{ij}}:

0.0088 0.8626 2.5355 1.0215 0.1496
0.0002 0.0178 0.0523 0.0211 0.0031

Value~of~\chi^2=\sum_{i=1}^2\sum_{j=1}^3\frac{(O_{ij}-E_{ij})^2}{E_{ij}}=4.67 \\\\ P-value=P(\chi^2>4.67|\chi^3\sim \chi^2_3)=0.1976\\\\ (R~code:~round(1-pchisq(4.67,3),4))

Since P-value>0.01, Option: Fail to reject the null hypothesis.

Conclusion: There is not sufficient evidence to support the claim that the proportion of people within each treatment group who experienced dizziness are different.